# Risk-Averse Constrained Reinforcement Learning with Optimized Certainty Equivalents

### v1

@@ 4.2 Results
The performance in the navigation tasks shows us that our method can strictly prevent any constraint violations, whether quantified continuously or not. In fact, it is the only method achieving strictly zero violations compared to other PPO-based constraint learning algorithms (see Table 5 in [19]). In some cases, constraint satisfaction also leads to higher rewards than the unconstrained vanilla agent (e.g., in the Push environment). This shows that constraints are not arbitrary but can guide the agent toward optimal task behavior.
@@ 5 Conclusion
An open problem resulting from this work asks whether or not full strong duality holds unconditionally. Also, while our framework is inherently risk-aware, it is more computationally intensive compared to risk-neutral methods (as each update of $\lambda$ and $t$ variables requires solving for an approximately optimal policy), as is also true for other risk-aware methods, e.g., [7, 12].
@@ A.1 Robustness in Value and Time
The so-called reward-based formulation in (4), as coined by Bisi et al. [6], Bonetti et al. [7], captures per-stage risk, in contrast to the return-based approach (3), which captures aggregate discounted risk. To see this, first note that the occupation measure can be reexpressed as
@@
We need to prove that it is valid to move the supremum over each $t_{i}$ to the objective. The proof follows that of Chow et al. [12].
@@
Paternain et al. [33, Theorem 1] proved strong duality holds for problems of the form
@@
which involve the expectation of the discounted sums of some bounded reward functions $r_{i}$, for $i\in\{0\}\cup[m]$. Thus, the problem (9) is exactly in the form of Paternain et al. [33], with the reward functions $r^{\prime}_{i}(s,a,t_{i})=t_{i}-\frac{1}{\beta}(t_{i}-r_{i}(s,a))_{+}$, for $i\in\{0\}\cup[m]$. Since the original $r_{i}$ are bounded and $t_{i}$ are fixed, we have that $r^{\prime}_{i}$ are bounded and we can readily apply Paternain et al. [33, Theorem 1]. ∎
@@
We assume that $\pi_{\theta}$, for $\theta\in\Theta\subset\mathbb{R}^{p}$ (with $\Theta$ some compact set), is an $\epsilon$-universal parametrization of measures in $\mathcal{P}(\mathcal{S})$, according to Paternain et al. [33, Definition 1] and that the parametrized-policy version of (7) is feasible. We proceed to show that there is almost no price to pay for the policy parametrization, in terms of duality gap.
Indeed, assuming that the parametrized optimization problem is feasible, we can utilize Paternain et al. [33, Theorem 2] and Assumption 3.4 to deduce that, for all $t\in\mathcal{I}$, the following inequalities hold:
@@
Thus, under the mere assumption that the sample gradient function $\widehat{\nabla}_{t,\lambda}\widehat{\mathcal{L}}(\pi_{\theta},t,\lambda)$ is $\widehat{L}-$Lipschitz continuous with respect to $\theta$ (which can be enforced, if necessary, by appropriate smoothing of the utility function $g$; a general smoothing strategy for generating smooth OCEs from non-smooth utilities can be found in Kouri and Surowiec [22, Example 2]), we obtain that
@@ A.6 Lipschitzness of the Lagrangian
    The first scenario, which is somewhat restrictive, relies on the assumption that, for each $(t,\lambda)\in\mathcal{T}\times\Lambda$, the maximization problem, with respect to $\theta$, admits a unique solution. In that case, we can readily utilize Lemma 2.2 of Shapiro [40], which states that $\mathcal{L}(\pi_{\theta^{*}(t,\lambda)},t,\lambda)$ is twice-continuously differentiable. Since $\mathcal{T}\times\Lambda$ is a compact set, this immediately implies that the function is $\ell$-smooth over $\mathcal{T}\times\Lambda$, for some constant $\ell>0$.
@@
    We note that the previous holds in several cases (typically with uniform exponent $\eta$). For example, this is true when $\max_{\theta\in\Theta}\mathcal{L}(\pi_{\theta},\cdot,\cdot)$ is sub-analytic (which, for example, is implied if $\mathcal{L}(\pi_{\theta},t,\lambda)$ is analytic, noting that this can easily be enforced our setting). Since $\mathcal{T}\times\Lambda$ is a compact and convex set, we can readily utilize Theorem 11 from [16] to deduce that under this general assumption, the function $\mathcal{L}(\pi_{\theta^{*}(t,\lambda)},t,\lambda)$ is $\ell$-smooth over $\mathcal{T}\times\Lambda$, irrespectively of the selections $\theta^{*}(t,\lambda)\in\arg\max_{\theta\in\Theta}\mathcal{L}(\pi_{\theta},t,\lambda)$.
@@
    Under certain qualification conditions, laid out in Section 4 of Shapiro [40] (that do not require the second-order strong sufficient conditions for the maximization problem over $\theta$), $\ell$-smoothness of $\max_{\theta\in\Theta}\mathcal{L}(\pi_{\theta},t,\lambda)$ is also guaranteed. We refer the reader to Shapiro [40] for additional details, since this analysis directly applicable in our setting.
@@
Then, we can substitute the original Lagrangian with the smoothed Lagrangian given in (23) and run Algorithm 1 on the surrogate problem. Following standard zeroth-order optimization techniques (e.g., see Pougkakiotis and Kalogerias [34]), we note that, in this case, the only algorithmic adjustments that need to be made relate to the computation of the sample gradients of $f$, which would require two (instead of one) evaluations of the policy oracle. In light of this discussion, we conclude that Assumption 3.8 can be utilized (almost) without loss of generality.
@@ A.8 Proof of Main Theorem
In this section, we provide the proof of Theorem 3.12. To prove the main theorem we will need the following auxiliary lemmas. The proof extends the analysis given in [25], since the presented algorithm only assumes having access to biased and inexact (sub)gradient samples.
@@
$\leq 2\left\lVert\nabla f(t,\lambda)\right\rVert^{2}+2(\sigma^{2}+\delta^{2})\,.$ (by Lemma A.2 of [25] with $M=1$ and Assumptions 3.6, 3.10)
@@
Taking expectations of both sides and using Lemma A.2 of [25], that $f(\cdot,\lambda)$ is $C$-Lipschitz (cf. Lemma 3.7), and Assumption 3.6,
@@
while, from [25, Enq. (23), (24)], we obtain
@@
From Lemma A.2, [25, Lemma A.2], and since $\mathrm{diam}(\Lambda)$ is finite, we get
@@
Following the steps in [25, Lemma D.4], using that $f(\cdot,\lambda)$ is $C$-Lipschitz (by Lemma 3.7) and Lemma A.2, we have
@@ B.1 Simulations
Constraint tasks are a central class of real-world RL problems with applications in robotics, autonomous vehicles, and industrial control. To simulate realistic conditions, we use the locomotion tasks from the Safety-Gymnasium suite [19]. Further details, including visualizations of both constrained and unconstrained agent behaviors, are available on their website [^5] https://safety-gymnasium.readthedocs.io/en/latest/environments/safe_velocity.html.
@@
where the velocity thresholds $\operatorname{vel}_{\text{threshold}}$ for the tested environments are listed in Table 3. As reported by Towers et al. [47], these thresholds are set to 50% of the maximum velocity achieved by each agent after PPO training for $10^{7}$ steps. At each time step, the agent’s instantaneous velocity is computed as:
@@ B.2 Proximal Policy Optimization – solver
We use Proximal Policy Optimization [39] as a solver to learn a policy (i.e., line 4 in Algorithm 1). Our method serves as a wrapper, described next, that modifies the agent’s raw reward to incorporate risk measures and constraints.
@@
We employ Generalized Advantage Estimation (GAE) [38] to estimate advantages in PPO. The hyperparameters used by the PPO agent is provided in Table 4.

### v2

@@ 4.2 Results
The performance in the navigation tasks shows us that our method can strictly prevent any constraint violations, whether quantified continuously or not. In fact, it is the only method achieving strictly zero violations compared to other PPO-based constraint learning algorithms (see Table 5 in [safety_gym]). In some cases, constraint satisfaction also leads to higher rewards than the unconstrained vanilla agent (e.g., in the Push environment). This shows that constraints are not arbitrary but can guide the agent toward optimal task behavior.
@@ 5 Conclusion
An open problem resulting from this work asks whether or not full strong duality holds unconditionally. Also, while our framework is inherently risk-aware, it is more computationally intensive compared to risk-neutral methods (as each update of $\lambda$ and $t$ variables requires solving for an approximately optimal policy), as is also true for other risk-aware methods, e.g., [bonetti_risk_averse_rl_coherent_risk, risk_contr_rl_percentile_risk].
@@ A.1 Robustness in Value and Time
The so-called reward-based formulation in (4), as coined by ijcai2020p632, bonetti_risk_averse_rl_coherent_risk, captures per-stage risk, in contrast to the return-based approach (3), which captures aggregate discounted risk. To see this, first note that the occupation measure can be reexpressed as
@@
We need to prove that it is valid to move the supremum over each $t_{i}$ to the objective. The proof follows that of risk_contr_rl_percentile_risk.
@@
constrained_rl_zero_duality_gap proved strong duality holds for problems of the form
@@
which involve the expectation of the discounted sums of some bounded reward functions $r_{i}$, for $i\in\{0\}\cup[m]$. Thus, the problem (9) is exactly in the form of constrained_rl_zero_duality_gap, with the reward functions $r^{\prime}_{i}(s,a,t_{i})=t_{i}-\frac{1}{\beta}(t_{i}-r_{i}(s,a))_{+}$, for $i\in\{0\}\cup[m]$. Since the original $r_{i}$ are bounded and $t_{i}$ are fixed, we have that $r^{\prime}_{i}$ are bounded and we can readily apply constrained_rl_zero_duality_gap. ∎
@@
We assume that $\pi_{\theta}$, for $\theta\in\Theta\subset\mathbb{R}^{p}$ (with $\Theta$ some compact set), is an $\epsilon$-universal parametrization of measures in $\mathcal{P}(\mathcal{S})$, according to constrained_rl_zero_duality_gap and that the parametrized-policy version of (7) is feasible. We proceed to show that there is almost no price to pay for the policy parametrization, in terms of duality gap.
Indeed, assuming that the parametrized optimization problem is feasible, we can utilize constrained_rl_zero_duality_gap and Assumption 3.4 to deduce that, for all $t\in\mathcal{I}$, the following inequalities hold:
@@
Thus, under the mere assumption that the sample gradient function $\widehat{\nabla}_{t,\lambda}\widehat{\mathcal{L}}(\pi_{\theta},t,\lambda)$ is $\widehat{L}-$Lipschitz continuous with respect to $\theta$ (which can be enforced, if necessary, by appropriate smoothing of the utility function $g$; a general smoothing strategy for generating smooth OCEs from non-smooth utilities can be found in epi_regularization_risk_meas_20), we obtain that
@@ A.6 Lipschitzness of the Lagrangian
    The first scenario, which is somewhat restrictive, relies on the assumption that, for each $(t,\lambda)\in\mathcal{T}\times\Lambda$, the maximization problem, with respect to $\theta$, admits a unique solution. In that case, we can readily utilize Lemma 2.2 of Shapiro_Extremal_functions, which states that $\mathcal{L}(\pi_{\theta^{*}(t,\lambda)},t,\lambda)$ is twice-continuously differentiable. Since $\mathcal{T}\times\Lambda$ is a compact set, this immediately implies that the function is $\ell$-smooth over $\mathcal{T}\times\Lambda$, for some constant $\ell>0$.
@@
    We note that the previous holds in several cases (typically with uniform exponent $\eta$). For example, this is true when $\max_{\theta\in\Theta}\mathcal{L}(\pi_{\theta},\cdot,\cdot)$ is sub-analytic (which, for example, is implied if $\mathcal{L}(\pi_{\theta},t,\lambda)$ is analytic, noting that this can easily be enforced our setting). Since $\mathcal{T}\times\Lambda$ is a compact and convex set, we can readily utilize Theorem 11 from [Hashmi_etal] to deduce that under this general assumption, the function $\mathcal{L}(\pi_{\theta^{*}(t,\lambda)},t,\lambda)$ is $\ell$-smooth over $\mathcal{T}\times\Lambda$, irrespectively of the selections $\theta^{*}(t,\lambda)\in\arg\max_{\theta\in\Theta}\mathcal{L}(\pi_{\theta},t,\lambda)$.
@@
    Under certain qualification conditions, laid out in Section 4 of Shapiro_Extremal_functions (that do not require the second-order strong sufficient conditions for the maximization problem over $\theta$), $\ell$-smoothness of $\max_{\theta\in\Theta}\mathcal{L}(\pi_{\theta},t,\lambda)$ is also guaranteed. We refer the reader to Shapiro_Extremal_functions for additional details, since this analysis directly applicable in our setting.
@@
Then, we can substitute the original Lagrangian with the smoothed Lagrangian given in (23) and run Algorithm 1 on the surrogate problem. Following standard zeroth-order optimization techniques (e.g., see PougkKal), we note that, in this case, the only algorithmic adjustments that need to be made relate to the computation of the sample gradients of $f$, which would require two (instead of one) evaluations of the policy oracle. In light of this discussion, we conclude that Assumption 3.8 can be utilized (almost) without loss of generality.
@@ A.8 Proof of Main Theorem
In this section, we provide the proof of Theorem 3.12. To prove the main theorem we will need the following auxiliary lemmas. The proof extends the analysis given in [pmlr-v119-lin20a], since the presented algorithm only assumes having access to biased and inexact (sub)gradient samples.
@@
$\leq 2\left\lVert\nabla f(t,\lambda)\right\rVert^{2}+2(\sigma^{2}+\delta^{2})\,.$ (by Lemma A.2 of [pmlr-v119-lin20a] with $M=1$ and Assumptions 3.6, 3.10)
@@
Taking expectations of both sides and using Lemma A.2 of [pmlr-v119-lin20a], that $f(\cdot,\lambda)$ is $C$-Lipschitz (cf. Lemma 3.7), and Assumption 3.6,
@@
while, from [pmlr-v119-lin20a, Enq. (23), (24)], we obtain
@@
From Lemma A.2, [pmlr-v119-lin20a, Lemma A.2], and since $\mathrm{diam}(\Lambda)$ is finite, we get
@@
Following the steps in [pmlr-v119-lin20a, Lemma D.4], using that $f(\cdot,\lambda)$ is $C$-Lipschitz (by Lemma 3.7) and Lemma A.2, we have
@@ B.1 Simulations
Constraint tasks are a central class of real-world RL problems with applications in robotics, autonomous vehicles, and industrial control. To simulate realistic conditions, we use the locomotion tasks from the Safety-Gymnasium suite [safety_gym]. Further details, including visualizations of both constrained and unconstrained agent behaviors, are available on their website [^5] https://safety-gymnasium.readthedocs.io/en/latest/environments/safe_velocity.html.
@@
where the velocity thresholds $\operatorname{vel}_{\text{threshold}}$ for the tested environments are listed in Table 3. As reported by gym, these thresholds are set to 50% of the maximum velocity achieved by each agent after PPO training for $10^{7}$ steps. At each time step, the agent’s instantaneous velocity is computed as:
@@ B.2 Proximal Policy Optimization – solver
We use Proximal Policy Optimization [ppo] as a solver to learn a policy (i.e., line 4 in Algorithm 1). Our method serves as a wrapper, described next, that modifies the agent’s raw reward to incorporate risk measures and constraints.
@@
We employ Generalized Advantage Estimation (GAE) [gae] to estimate advantages in PPO. The hyperparameters used by the PPO agent is provided in Table 4.